An engineer weighs a sample of mercury ( density= 13.6 × 103 kg/m3) and finds that the weight of the sample is 4.3 N.
What is the sample’s volume? The acceleration of gravity is 9.81 m/s2 .
Answer in units of m3.
1
Expert's answer
2011-01-25T09:20:00-0500
We have to find the mass of the sample:
m = F/g = 4.3/9.81 = 0.438 kg
Thus the volume of the sample is
V = m/ρ = 0.438/13.6 × 103 = 3.2 × 10-5 m3.